# PackUV: Packed Gaussian UV Maps for 4D Volumetric Video

### v1

@@ Abstract
[…]
@@ 4.2.2 Gaussian Labeling
Each Gaussian $g_{i}$ has mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{3}$, diagonal scales $\mathbf{s}_{i}\in\mathbb{R}^{3}$, and rotation $\mathbf{q}_{i}$ (unit quaternion). Its 3D covariance is represented using $\mathbf{R}(\cdot)$ is the rotation matrix and $\mathbf{S}(\mathbf{s})=\mathrm{diag}(\mathbf{s})$ [32].
@@
$\nabla_{\boldsymbol{\theta}_{i}}\mathcal{L}\;\leftarrow\;D_{i}\,\nabla_{\boldsymbol{\theta}_{i}}\mathcal{L}\,,\qquad D_{i}\in\{0,1\},$
where $\boldsymbol{\theta}_{i}$ includes position, scale, rotation, color, and opacity. We additionally reset optimizer momentum for static Gaussians periodically to prevent drift. When densifying, child Gaussians inherit the parent’s dynamic/static label to preserve the dynamic ratio across training.
@@ 12.1 Gaussian Labeling:
1:Gaussian parameters $\{\boldsymbol{\mu}_{i},\mathbf{s}_{i},\mathbf{q}_{i}\}_{i=1}^{N}$, camera views $\mathcal{C}$, flow masks $\{M^{c}\}_{c\in\mathcal{C}}$
@@
10:     $\boldsymbol{\mu}_{i,\text{cam}}\leftarrow\mathbf{T}_{c}\cdot\boldsymbol{\mu}_{i}$
@@
12:     Compute Jacobian: $\mathbf{J}=\begin{bmatrix}f_{x}/z&0&-f_{x}x/z^{2}\\ 0&f_{y}/z&-f_{y}y/z^{2}\end{bmatrix}$ at $\boldsymbol{\mu}_{i,\text{cam}}$
@@
16:     $\mathbf{m}_{i,c}\leftarrow\text{NDC2Pixel}(\mathbf{P}_{c}\cdot\boldsymbol{\mu}_{i})$ $\triangleright$ Project mean to pixels
@@
29:     for $\mathbf{p}\in\{\mathbf{m}_{i,c}+\boldsymbol{\delta}\mid\|\boldsymbol{\delta}\|_{\infty}\leq r\}$ do
@@ 13 Viewer
We also built a custom viewer for rendering our PackUV atlas maps on top of the Tiny Gaussian Splatting Viewer by Li Ma, which originally supports OpenGL and CUDA backends for static scenes. Our viewer extends this functionality to enable interactive camera control, real-time playback at arbitrary frame rates, and frame-by-frame navigation through dynamic sequences.
Rendering is performed using OpenGL shaders and shader storage buffer objects (SSBOs). Each video frame is represented as a PackUV atals stored as a npz file, decreasing our loading latency compared to .ply files. And to accelerate playback, we cache a flattened, OpenGL-ready version of each frame on the CPU. At runtime, when advancing to a new frame, the viewer copies this preprocessed buffer to the GPU for rendering. Between frame updates, the current buffer remains active, allowing continuous rendering and smooth scene exploration.
Excluding I/O, our renderer achieves >200>\!200 FPS; including I/O, it maintains $\geq\!30$ FPS on a NVIDIA 3090 GPU, with performance independent of sequence length due to buffering. We set the number of GS in the range of 300K-500K for every scene.
@@ 15 Limitations and Future Work
While our method achieves high-quality 4D neural video renderings, some limitations remain that requires further exploration. A primary challenge arises from the inherently unstructured nature of 3D Gaussian representations. Although our approach introduces structure through UV projection, it still requires enforcing a large spatial arrangement to capture the fine details of real-world scenes. Furthermore, while PackUV enables the use of video coding infrastructure by mapping to a single frame, discovering more optimal mappings could further improve storage efficiency. Another promising direction is to make this representation compatible with AR/VR devices, thereby enabling direct 4D streaming for immersive applications.

### v2

@@ Abstract
Volumetric videos offer immersive 4D experiences, but remain difficult to reconstruct, store, and stream at scale. Existing Gaussian Splatting based methods achieve high-quality reconstruction but break down on long sequences, temporal inconsistency, and fail under large motions and disocclusions. Moreover, their outputs are typically incompatible with conventional video coding pipelines, preventing practical applications. We introduce PackUV, a novel 4D Gaussian representation that maps all Gaussian attributes into a sequence of structured, multi-scale UV atlases, enabling compact, image-native storage. To fit this representation from multi-view videos, we propose PackUV-GS, a temporally consistent fitting method that directly optimizes Gaussian parameters in the UV domain. A flow-guided Gaussian labeling and video keyframing module identifies dynamic Gaussians, stabilizes static regions, and preserves temporal coherence even under large motions and disocclusions. The resulting UV atlas format is the first unified volumetric video representation compatible with standard video codecs (e.g., FFV1) without losing quality, enabling efficient streaming within existing multimedia infrastructure. To evaluate long-duration volumetric capture, we present PackUV-2B, the largest multi-view 4D dataset to date, featuring more than 50 synchronized cameras, substantial motion, and frequent disocclusions across 100 sequences and 2B (billion) frames. Extensive experiments demonstrate that our method surpasses existing baselines in rendering fidelity while scaling to sequences up to 30 minutes with consistent quality.
@@ 4.2.2 Gaussian Labeling
Each Gaussian $g_{i}$ has mean $\bm{\mu}_{i}\in\mathbb{R}^{3}$, diagonal scales $\mathbf{s}_{i}\in\mathbb{R}^{3}$, and rotation $\mathbf{q}_{i}$ (unit quaternion). Its 3D covariance is represented using $\mathbf{R}(\cdot)$ is the rotation matrix and $\mathbf{S}(\mathbf{s})=\mathrm{diag}(\mathbf{s})$ [32].
@@
$\nabla_{\bm{\theta}_{i}}\mathcal{L}\;\leftarrow\;D_{i}\,\nabla_{\bm{\theta}_{i}}\mathcal{L}\,,\qquad D_{i}\in\{0,1\},$
where $\bm{\theta}_{i}$ includes position, scale, rotation, color, and opacity. We additionally reset optimizer momentum for static Gaussians periodically to prevent drift. When densifying, child Gaussians inherit the parent’s dynamic/static label to preserve the dynamic ratio across training.
@@ 7 Conclusion
ACKNOWLEDGEMENTS
This research was supported by ONR DURIP grant N00014-23-1-2804, NSF CAREER award #\#2143576, and an Amazon Cloud Credits award. We also want to thank Emre Arslan for building the first version of the viewer. We also appreciate the time and efforts of all the participants for data collection.
@@ 12.1 Gaussian Labeling:
1:Gaussian parameters $\{\bm{\mu}_{i},\mathbf{s}_{i},\mathbf{q}_{i}\}_{i=1}^{N}$, camera views $\mathcal{C}$, flow masks $\{M^{c}\}_{c\in\mathcal{C}}$
@@
10:     $\bm{\mu}_{i,\text{cam}}\leftarrow\mathbf{T}_{c}\cdot\bm{\mu}_{i}$
@@
12:     Compute Jacobian: $\mathbf{J}=\begin{bmatrix}f_{x}/z&0&-f_{x}x/z^{2}\\ 0&f_{y}/z&-f_{y}y/z^{2}\end{bmatrix}$ at $\bm{\mu}_{i,\text{cam}}$
@@
16:     $\mathbf{m}_{i,c}\leftarrow\text{NDC2Pixel}(\mathbf{P}_{c}\cdot\bm{\mu}_{i})$ $\triangleright$ Project mean to pixels
@@
29:     for $\mathbf{p}\in\{\mathbf{m}_{i,c}+\bm{\delta}\mid\|\bm{\delta}\|_{\infty}\leq r\}$ do
@@ 13 Viewer
The viewer renders a dynamic 4D Gaussian scene as a time-ordered sequence of Gaussian frames. At time step $t_{k}$, the scene is represented as
$\mathcal{G}(t_{k})=\{g_{i}^{(k)}\}_{i=1}^{N_{k}},$
where each Gaussian $g_{i}^{(k)}$ stores a 3D mean $\bm{\mu}_{i}$, covariance $\bm{\Sigma}_{i}$, opacity $\alpha_{i}$, and appearance parameters (e.g., SH coefficients). A stream manager loads PackUV-encoded frame data sequentially, maintains a small look-ahead cache, and exposes both CPU-side Gaussian data and GPU-ready packed metadata for the current frame.
During playback, the active frame index is advanced according to the target frame rate,
$k(t)=\lfloor ft\rfloor\bmod N,$
where $f$ is the playback FPS and $N$ is the number of frames. When a new frame is selected, its packed representation is uploaded to the renderer and the Gaussian set is updated before drawing.
Rendering follows the standard 3D Gaussian splatting pipeline. A Gaussian in world coordinates is projected into screen space using the current camera pose and intrinsics. For a 3D point $\bm{X}$, projection is given by
$\bm{X}_{c}=\bm{R}\bm{X}+\bm{t},\hskip 28.80008pt\tilde{\bm{x}}=\bm{K}\bm{X}_{c},$
with image coordinates obtained by perspective division. The projected 2D covariance is approximated as
$\bm{\Sigma}_{2D}=\bm{J}\bm{W}\bm{\Sigma}_{3D}\bm{W}^{\top}\bm{J}^{\top},$
where $\bm{W}$ is the local world-to-camera transform and $\bm{J}$ is the Jacobian of the projection function. The resulting ellipse defines the screen-space footprint of each splat.
To support correct transparency, Gaussians are depth-sorted with respect to the current camera and composited back-to-front:
$\bm{C}=\sum_{i=1}^{M}T_{i}\alpha_{i}\bm{c}_{i},\hskip 28.80008ptT_{i}=\prod_{j<i}(1-\alpha_{j}),$
where $\bm{c}_{i}$ is the color contribution of the $i$-th splat. The viewer additionally supports multiple shading modes, including depth visualization and spherical harmonics shading, making it suitable for interactive inspection of dynamic PackUV-based 4DGS sequences.
Excluding I/O, our viewer achieves >200>\!200 FPS; including I/O, it maintains $\geq\!30$ FPS on a NVIDIA 4090 GPU with a set number of GS in the range of 300K-500K for every scene. The viewer’s rendering performance independent of sequence length due to buffering.
@@ 15 Limitations and Future Work
While our method achieves high-quality 4D neural video renderings, some limitations remain that requires further exploration. A primary challenge arises from the inherently unstructured nature of 3D Gaussian representations. Although our approach introduces structure through UV projection, it still requires enforcing a large spatial arrangement to capture the fine details of real-world scenes.
PackUV can sometimes produce dragged Gaussian artifacts when the optical flow is estimated inaccurately, which becomes a bottleneck in the pipeline. Although setting more keyframes and reducing optical flow threshold could fix the problem, improving the flow prediction module or adopting a more robust mapping strategy could lead to better temporal consistency.
Furthermore, while PackUV enables the use of video coding infrastructure by mapping to a single frame, discovering more optimal mappings could further improve storage efficiency. Another promising direction is to make this representation compatible with AR/VR devices, thereby enabling direct 4D streaming for immersive applications.